You are planning to buy a house, you can either lease the house from the

society or purchase it with a six-year loan. The house you wish to buy costs Rs.30,000,000. The

dealer has a special leasing arrangement where you pay Rs.1,700,000 today and Rs.450,000 per

month for the next six years. If you purchase the house, you will pay it off in monthly payments

over the next six years at required rate of return of 6 percent per annum. You believe that you

will be able to sell the house for Rs.35,000,000 in six years.

Required to calculate and answer the following :

1. What option will be beneficial for you either buying house in single payment or buying

on lease.

2. Assume you purchase the house in onetime payment and sold it for 35,000,000 after six

years on same interest rate compounding monthly will this be a good and profitable deal

for you.


1
Expert's answer
2021-01-16T18:14:43-0500

1)Let use the financial formulas of the Excel. Let's calculate Net Present Value (NPV) for both cases:

"NPV=P_0+PV(6\%/12;6\cdot12;450000;)=-1700000-27 152 781.27=-28 852 781."

The value listed above is NPV for the case of lease the house from the society.

At the same time for buiyng in single payment "NPV=-30000000". So, the benefit comparing the case of the lease is "1 147 218.73" which makes lease preferable.

2) Selling the house after six years will give us the profit of


"PV(6\%/12;6\cdot12;;-35000000)=24 440 585"

In case of lease "NPV" will be -28 852 781+24 440 585= -4412196.

In case of buiyng in single payment NPV will be -30000000+24 440 585=-5559414.

In this case the lease is also preferable.